is these true or false ?and explain why

a)if f(n)=O(g(n)) then 2^(f(n)=O(2^(g(n)))... please solve without lim

b)if f(n)=o(g(n)) then 2^(f(n)=o(2^(g(n)))... please solve without lim
Answer #1

FALSE

If f(n) = O(g(n)),
2^(f(n)) not equal to O(2^g(n)))

Let, f(n) = 2log n and g(n) = log n
(Assume log is to the base 2)

We know, 2log n <= c(log n) therefore f(n) = O(g(n))

2^(f(n)) = 2^log n^2 = n^2
2^(g(n)) = 2^log n = n

We know that
n^2 is not O(n)

Therefore, 2^(f(n)) not equal to O(2^g(n)))